# Branch, or Layer? Zeroth-Order Optimization for Continual Learning of Vision-Language Models

### v2

@@ Introduction
Continual Learning (CL) has witnessed significant advancements in convolutional architectures (e.g., ResNet (erd; res_2; pose) and ViT(vit_1; vit_2; vit_3)). Recently, Vision-Language Models-based Continual Learning (VLCL) approaches have attracted growing research attention. Particularly, CLIP-based methods (clip_cl2; lwf-vr; zhao2; mod-x) have demonstrated robust continual learning capabilities. However, these methods require full-parameter fine-tuning of CLIP models, which incurs substantial computational overhead. To overcome this critical bottleneck, Parameter-Efficient Fine-Tuning (PEFT) strategies (peft; peft1; peft2; dmnsp) have recently emerged as a compelling alternative. These techniques make it possible to achieve competitive CL performance with significantly reduced resource consumption. For instance, the VLCL method MoE4Adapter (MoE4Adapters) leverages a PEFT approach to effectively address this limitation.
Existing VLCL methods predominantly employ First-Order (FO) optimization strategy (overview_opti), which update parameters using precise gradients derived from backpropagation. While valued for their stable directional guidance, this approach becomes a drawback in the context of PEFT. Its deterministic update paths limit exploration during training , and the low-dimensional subspace to which PEFT confines optimization makes these methods susceptible to converging to sharp local optima that overfit the current task (SharpMinima; sam), leading to a performance drop when faced with new tasks and potentially exacerbating catastrophic forgetting. To explore solutions with stronger generalization capabilities, Zeroth-Order (ZO) optimization offers a promising alternative (zeroflow). Unlike traditional FO optimization, this method forgoes precise gradients from backpropagation, instead estimating performance with random perturbations, making it less prone to getting trapped in local minima when exploring a constrained space (mezo; maskzo), hence potentially applied in PEFT-based VLCL.
There is already a vast amount of ZO finetuning researches (mezo; maskzo), while they lacks of considering its common strategy of fully replacing FO optimization is applicable to VLCL, given the documented optimization disparities between modality branches (vlgap; clap4clip; vldis1; vldis2; vldis3) and the inherent sensitivity of ZO’s perturbation-based approach. In this paper, we initiate the study of applying ZO optimization into VLCL, and aim to answer the following key question: How can ZO be integrated in VLCL settings, and can it boost overall performance?
@@ A Preliminary for ZO Optimization in PEFT-based VLCL
Traditional optimization methods in continual learning primarily rely on FO gradient descent (fo_1; fo_2), updating model parameters $\theta$ based on precise gradients computed via backpropagation:
@@
where $\Delta$ is a random directional vector, and $\varepsilon$ is a small perturbation scale. ZO methods introduce gradient stochasticity (derivative), which may improve the ability to explore non-convex loss landscapes and escape poor local minima. However, the reliance on randomized perturbations can lead to variance in gradient estimates, whose effect may vary across different model architectures or modalities. To validate this point, we first establish a baseline and investigate the most straightforward strategy: a complete replacement of the FO optimizer with ZO in PEFT-based VLCL, which means adopt ZO in both vision-language branches and overall trainable units. However, our subsequent experiments demonstrate this strategy leads to severe training instability, evidenced by loss oscillations during the convergence process, and results in a significant performance degradation.
Recent research suggests that partially incorporating exploratory while high-variance ZO estimators into the model architecture improves its global optimization performance (hybridzo1; hybridzo2). We translate this advance to the context of PEFT-based VLCL. As shown in Figure 1, we adopt a empirical exploration to progressively investigate how to best apply the ZO optimization in VLCL. To explore suitable application paradigms, we designed a comprehensive set of experiments applying ZO at varying levels of granularity, from entire modality branches to fine-grained partially trainable units. However, given the vast diversity of configurations arising from combinations of modality branches and PEFT trainable units, constructing a single, unified theoretical framework is intractable. Consequently, we conduct a comprehensive empirical analysis to investigate both the differential impacts of applying ZO across modality branches and the distinct characteristics of employing it in various trainable layer configurations (such as continuous and interleaved layers), which allows us to achieve optimal results.
@@
Baseline. We choose MoE4Adapter (MoE4Adapters) as the SOTA baseline, which incorporates Mixture of Experts (MoE) into CLIP for VLCL. To explore PEFT alternatives, we also replaces the MoE modules with Low-Rank Adaptation (LoRA (lora)) modules. For LoRA results, we present a part of them and put the remaining in supplementary material.
Implementation details. All experiments employ the CLIP-ViT-B/16 backbone architecture. Following baseline’s paradigm, the CLIP backbone remains frozen, with only task-specific adapters (MoE or LoRA modules) being trainable. For the ZO-based method, we adopt a more conservative ZO strategy which evaluates multiple candidate updates and selectively applies the one that yields the lowest loss(zeroflow). Hyperparameters including perturbation scale $\varepsilon=0.001$ for ZO gradients and FO/ZO mixing ratio $\lambda=1$, are validated on the first task and retained for subsequent tasks.
@@
Table 1: How ZO optimization affects VLCL through different branches (CLIP). w/ ZO denotes the branch (Dual, Vision, or Language) where ZO optimization is applied. The $\dagger$} indicates that the MoE modules in baseline are replaced with LoRA.
@@
MoE4Adapter 80.47 86.97 77.52 85.21 52.13 60.55 65.36 71.53
    Dual w/ ZO 69.29 77.36 67.64 75.88 42.40 47.64 58.56 65.92
    Vision w/ ZO 76.05 83.93 72.98 82.08 49.65 57.90 62.54 69.84
    Language w/ ZO 80.94 87.00 76.74 85.03 49.14 58.69 64.38 70.38
MoE4Adapter† 80.44 87.10 79.66 86.34 51.93 59.80 64.34 71.79
    Dual w/ ZO 71.07 78.37 69.70 76.82 44.53 52.71 58.71 65.97
    Vision w/ ZO 75.86 83.86 73.73 82.25 49.98 58.11 62.70 69.88
    Language w/ ZO 77.47 85.40 79.63 87.01 49.90 58.92 64.33 70.59
@@
Figure 2: How ZO optimization affects loss convergence of VLCL across different branches (CLIP). w/ ZO denotes the branch (Dual, Vision, or Language) where ZO optimization is applied.
@@
Building on these observations, the effectiveness of coarse-grained integration motivates us to explore the potential of synergistic optimization between ZO and FO methods. To further investigate this synergy, we consider a fine-grained perspective: Can layer-wise allocation of ZO and FO optimization within modality-specific enhance CL performance?
@@
(a) Du. w/Hop-odd.
(b) Vis. w/Hop-odd.
(c) Lan. w/Hop-odd.
Figure 3: Analyzing convergence behavior of VLCL in Hop-odd across Dual (Du.), Vision (Vis.), Language (Lan.).
@@ A New Enhancement from Vision Discrepancy
Figure 8: Significance analysis of performance across Dual, Vision, Language.
@@
Continual learning of VLM. Recent advances in CL have witnessed the emergence of VLMs as promising solutions to mitigate catastrophic forgetting through their generalized multimodal representations. Pioneering studies (lwf-vr; clip_cl2) demonstrate that pretrained VLMs like CLIP (clip) inherently possess remarkable continual learning capabilities even without fine-tuning. PROOF (lwf-vr) further enhances CL robustness by integrating multimodal cues with adaptive mapping strategies. Nevertheless, conventional VLM-based approaches predominantly emphasize task-specific feature acquisition for new domains (ralf), inadvertently compromising the integrity of previously learned representations and leading to progressive performance degradation. To address this limitation, recent efforts adopt PEFT techniques (peft1; peft2) that selectively update lightweight modules while maintaining frozen backbone parameters. MoE4Adapters (MoE4Adapters) introduces a mixture-of-experts (MoE) architecture, enabling task-specific feature specialization without cross-task interference and achieving state-of-the-art performance. However, these methods universally rely on FO optimization, inherently restricting their capacity to explore optimal parameter trajectories during optimization.
Optimization for continual learning. From an optimization view, existing CL methods predominantly focus on reconciling the stability-plasticity dilemma (staplast) through gradient technology (unigrad). Orthogonal parameter updates (grad1; grad2; grad3; grad4) and sharpness-aware minimization (flat1; flat2; flat3) represent two mainstream directions, which respectively aim to decouple task-specific gradients and converge to flat loss minima for improved generalization. While these strategies enhance optimization stability, they often inadequately address the exploration-exploitation trade-off, as deterministic gradient descent trajectories tend to converge to suboptimal local minima with limited perturbation resilience. Emerging ZO optimization(zo1; derivative) presents a paradigm shift by employing gradient-free stochastic perturbations to estimate descent directions. This approach offers dual advantages: 1) inherent stochasticity facilitates escape from local optima through controlled parameter space exploration, and 2) elimination of backward propagation reduces memory overhead by avoiding gradient matrix storage.
Our work. This work systematically investigates how ZO optimization can be effectively adapted to VLCL, addressing catastrophic forgetting by harmonizing the inter-modal asymmetry of VLMs with the stochastic perturbations of ZO optimization.
@@ Conclusion
Acknowledgement
This work is supported in part by the National Natural Science Foundation of China (62192783), Young Elite Scientists Sponsorship Program by CAST (2023QNRC001), and China Scholarship Council.

### v3

@@ Introduction
Continual Learning (CL) has witnessed significant advancements in convolutional architectures (e.g., ResNet (Feng et al. 2022b; Rebuffi et al. 2017; Feng et al. 2022a) and ViT(Wang et al. 2022b, a; Gao et al. 2024)). Recently, Vision-Language Models-based Continual Learning (VLCL) approaches have attracted growing research attention. Particularly, CLIP-based methods (Thengane et al. 2022; Ding et al. 2022; Zhao et al. 2023; Ni et al. 2023) have demonstrated robust continual learning capabilities. However, these methods require full-parameter fine-tuning of CLIP models, which incurs substantial computational overhead. To overcome this critical bottleneck, Parameter-Efficient Fine-Tuning (PEFT) strategies (Zhang et al. 2025a; Houlsby et al. 2019; Hung et al. 2019; Kang et al. 2025) have recently emerged as a compelling alternative. These techniques make it possible to achieve competitive CL performance with significantly reduced resource consumption. For instance, the VLCL method MoE4Adapter (Yu et al. 2024) leverages a PEFT approach to address this limitation.
Existing VLCL methods predominantly employ First-Order (FO) optimization strategy (Ruder 2016), which update parameters using precise gradients derived from backpropagation. While valued for their stable directional guidance, this approach becomes a drawback in the context of PEFT. Its deterministic update paths limit exploration during training , and the low-dimensional subspace to which PEFT confines optimization makes these methods susceptible to converging to sharp local optima that overfit the current task (Keskar et al. 2017; Möllenhoff and Khan 2023), leading to a performance drop when faced with new tasks and potentially exacerbating catastrophic forgetting. To explore solutions with stronger generalization capabilities, Zeroth-Order (ZO) optimization offers a promising alternative (Feng et al. 2025). Unlike traditional FO optimization, this method forgoes precise gradients from backpropagation, instead estimating performance with random perturbations, making it less prone to getting trapped in local minima when exploring a constrained space (Malladi et al. 2023; Zhang et al. 2025b), hence potentially applied in PEFT-based VLCL.
There is already a vast amount of ZO finetuning researches (Malladi et al. 2023; Zhang et al. 2025b), while they lacks of considering its common strategy of fully replacing FO optimization is applicable to VLCL, given the documented optimization disparities between modality branches (Liang et al. 2022; Jha et al. 2024; Peng et al. 2022; Sun et al. 2020; Cheng et al. 2024) and the inherent sensitivity of ZO’s perturbation-based approach. In this paper, we initiate the study of applying ZO optimization into VLCL, and aim to answer the following key question: How can ZO be integrated in VLCL settings, and can it boost overall performance?
@@ A Preliminary for ZO Optimization in PEFT-based VLCL
Traditional optimization methods in continual learning primarily rely on FO gradient descent (Cha et al. 2020; Hadsell et al. 2020), updating model parameters $\theta$ based on precise gradients computed via backpropagation:
@@
where $\Delta$ is a random directional vector, and $\varepsilon$ is a small perturbation scale. ZO methods introduce gradient stochasticity (Berahas et al. 2022), which may improve the ability to explore non-convex loss landscapes and escape poor local minima. However, the reliance on randomized perturbations can lead to variance in gradient estimates, whose effect may vary across different model architectures or modalities. To validate this point, we first establish a baseline and investigate the most straightforward strategy: a complete replacement of the FO optimizer with ZO in PEFT-based VLCL, which means adopt ZO in both vision-language branches and overall trainable units. However, our subsequent experiments demonstrate this strategy leads to severe training instability, evidenced by loss oscillations during the convergence process, and results in a significant performance degradation.
Recent research suggests that partially incorporating exploratory while high-variance ZO estimators into the model architecture improves its global optimization performance (Talaei et al. 2025; Chen et al. 2025). We translate this advance to the context of PEFT-based VLCL. As shown in Figure 1, we adopt a empirical exploration to progressively investigate how to best apply the ZO optimization in VLCL. To explore suitable application paradigms, we designed a comprehensive set of experiments applying ZO at varying levels of granularity, from entire modality branches to fine-grained partially trainable units. However, given the vast diversity of configurations arising from combinations of modality branches and PEFT trainable units, constructing a single, unified theoretical framework is intractable. Consequently, we conduct a comprehensive empirical analysis to investigate both the differential impacts of applying ZO across modality branches and the distinct characteristics of employing it in various trainable layer configurations (such as continuous and interleaved layers), which allows us to achieve optimal results.
@@
Baseline. We choose MoE4Adapter (Yu et al. 2024) as the SOTA baseline, which incorporates Mixture of Experts (MoE) into CLIP for VLCL. To explore PEFT alternatives, we also replaces the MoE modules with Low-Rank Adaptation (LoRA (Hu et al. 2022)) modules. For LoRA results, we present a part of them and put the remaining in supplementary material.
Implementation details. All experiments employ the CLIP-ViT-B/16 backbone architecture. The CLIP backbone remains frozen, with only task-specific adapters (MoE or LoRA modules) being trainable. For the ZO-based method, we adopt a more conservative ZO strategy which evaluates multiple candidate updates and selectively applies the one that yields the lowest loss(Feng et al. 2025). Hyperparameters including perturbation scale $\varepsilon=0.001$ for ZO gradients and FO/ZO mixing ratio $\lambda=1$, are validated on the first task and retained for subsequent tasks.
@@
Baseline 80.47 86.97 77.52 85.21 52.13 60.55 65.36 71.53
Du. w/ ZO 69.29 77.36 67.64 75.88 42.40 47.64 58.56 65.92
Vis. w/ ZO 76.05 83.93 72.98 82.08 49.65 57.90 62.54 69.84
Lan. w/ ZO 80.94 87.00 76.74 85.03 49.14 58.69 64.38 70.38
Baseline† 80.44 87.10 79.66 86.34 51.93 59.80 64.34 71.79
Du. w/ ZO 71.07 78.37 69.70 76.82 44.53 52.71 58.71 65.97
Vis. w/ ZO 75.86 83.86 73.73 82.25 49.98 58.11 62.70 69.88
Lan. w/ ZO 77.47 85.40 79.63 87.01 49.90 58.92 64.33 70.59
Table 1: How ZO optimization affects VLCL in different branches (CLIP). w/ ZO denotes the branch (Du. (Dual), Vis. (Vision), or Lan. (Language)) where ZO optimization is applied. The $\dagger$} indicates MoE modules in baseline are replaced with LoRA.
@@
Figure 2: How ZO optimization affects loss convergence of VLCL in different branches (CLIP). w/ ZO denotes the branch (Dual, Vision, or Language) where ZO applied.
@@
Building on these observations, the effectiveness of coarse-grained integration motivates us to explore the potential of synergistic optimization between ZO and FO methods. We further consider a fine-grained perspective: Can layer-wise allocation of ZO and FO optimization within modality-specific enhance CL performance?
@@
(a) Du. w/Hop-odd.
(b) Lan. w/Hop-odd.
Figure 3: Analyzing convergence behavior of VLCL in Hop-odd across Dual (Du.) and Language (Lan.) branches.
@@ A New Enhancement from Vision Discrepancy
Figure 8: Significance analysis of performance across Dual (Du.), Vision (Vis.), Language (Lan.).
@@
Continual learning of VLM. Recent advances in CL have witnessed the emergence of VLMs as promising solutions to mitigate catastrophic forgetting through their generalized multimodal representations. Pioneering studies (Ding et al. 2022; Thengane et al. 2022) demonstrate that pretrained VLMs like CLIP (Radford et al. 2021) inherently possess remarkable continual learning capabilities even without fine-tuning. PROOF (Ding et al. 2022) further enhances CL robustness by integrating multimodal cues with adaptive mapping strategies. Nevertheless, conventional VLM-based approaches predominantly emphasize task-specific feature acquisition for new domains (Huang et al. 2024), inadvertently compromising the integrity of previously learned representations and leading to progressive performance degradation. To address this limitation, recent efforts adopt PEFT techniques (Houlsby et al. 2019; Hung et al. 2019) that selectively update lightweight modules while maintaining frozen backbone parameters. MoE4Adapters (Yu et al. 2024) introduces a mixture-of-experts (MoE) architecture, enabling task-specific feature specialization without cross-task interference and achieving state-of-the-art performance. However, these methods universally rely on FO optimization, inherently restricting their capacity to explore optimal parameter trajectories during optimization.
Optimization for continual learning. From an optimization view, existing CL methods predominantly focus on reconciling the stability-plasticity dilemma (Lu et al. 2025) through gradient technology (Li et al. 2024). Orthogonal parameter updates (Saha et al. 2021; Lin et al. 2022; Lopez-Paz and Ranzato 2017; Farajtabar et al. 2020) and sharpness-aware minimization (Deng et al. 2021; Shi et al. 2021; Bian et al. 2024) represent two mainstream directions, which respectively aim to decouple task-specific gradients and converge to flat loss minima for improved generalization. While these strategies enhance optimization stability, they often inadequately address the exploration-exploitation trade-off, as deterministic gradient descent trajectories tend to converge to suboptimal local minima with limited perturbation resilience. Emerging ZO optimization(Nesterov and Spokoiny 2017; Berahas et al. 2022) presents a paradigm shift by employing gradient-free stochastic perturbations to estimate descent directions. This approach offers dual advantages: 1) inherent stochasticity facilitates escape from local optima through controlled parameter space exploration, and 2) elimination of backward propagation reduces memory overhead by avoiding gradient matrix storage.
Our work. This work investigates how ZO optimization can be effectively adapted to VLCL, addressing catastrophic forgetting by harmonizing the inter-modal asymmetry of VLMs with the stochastic perturbations of ZO optimization.
@@ Conclusion
Acknowledgements
This work is supported in part by the National Natural Science Foundation of China (62192783), Young Elite Scientists Sponsorship Program by CAST (2023QNRC001), and China Scholarship Council with award number 202406240114.